# Invariance of deep image quality metrics to affine transformations

### v2

@@ Abstract
Deep architectures are the current state-of-the-art in predicting subjective image quality. Usually, these models are evaluated according to their ability to correlate with human opinion in databases with a range of distortions that may appear in digital media. However, these oversee affine transformations which may represent better the changes in the images actually happening in natural conditions. Humans can be particularly invariant to these natural transformations, as opposed to the digital ones.
In this work, we evaluate state-of-the-art deep image quality metrics by assessing their invariance to affine transformations, specifically: rotation, translation, scaling, and changes in spectral illumination. Here invariance of a metric refers to the fact that certain distances should be neglected (considered to be zero) if their values are below a threshold. This is what we call invisibility threshold of a metric. We propose a methodology to assign such invisibility thresholds for any perceptual metric. This methodology involves transformations to a distance space common to any metric, and psychophysical measurements of thresholds in this common space. By doing so, we allow the analyzed metrics to be directly comparable with actual human thresholds. We find that none of the state-of-the-art metrics shows human-like results under this strong test based on invisibility thresholds.
This means that tuning the models exclusively to predict the visibility of generic distortions may disregard other properties of human vision as for instance invariances or invisibility thresholds. The data and code is publicly available to test other metrics [^1] https://github.com/Rietta5/InvarianceTestIQA.
Index Terms:
Perceptual Metrics, Perception Thresholds, Invariance to Affine Transformations
I Introduction
Deep architectures are the current state-of-the-art in predicting subjective image quality. These models, usually referred to as perceptual metrics, are often used as measures to optimize other image processing tasks [1, 2] so assessing their performance is critical. Usually, these models are evaluated (or even tuned) according to their ability to correlate with human opinion in databases including a wide range of generic distortions [3, 4]. In fundamental terms, this means trying to predict the visibility of generic distortions. However, focus on generic distortions may be a problem to take into account other relevant phenomena [5], and human vision may also be described in terms of perceptual constancies or invariances [6, 7, 8, 9, 10, 11].
@@ I Introduction
In fact, the spirit of the influential SSIM was focused on measuring changes of structure [13] to achieve invariances to irrelevant transformations [14]. Moreover, Wang and Simoncelli [12] decomposed generic distortions into structural and non-structural components so that the part not affecting the structure (e.g. affine transformations) could be processed, and weighted, differently. On the other hand, metrics with bio-inspired, explainable architecture [15, 16, 17, 18, 19, 20, 5, 4] work in multi-scale / multi-orientation representations where invariances can be introduced by means of appropriate poolings as in the representations [21]. This sort of poolings are thought to happen in the visual brain leading to invariances and texture metamers [22].
@@
In this work, we compare the ability of metrics to be invariant to affine transformations in the same way as humans are. In particular, we propose to evaluate the metrics from the point of view of human detection thresholds: by assessing the invariance of the metrics to image transformations that are irrelevant (or invisible) to human observers. For example, the classical literature on visual thresholds determines the intensity of certain affine transformations which is invisible for humans [24, 25, 26, 27]. The sizes of invisibility thresholds are related to the more general concept of invariance to transformations. By definition, transformations whose intensity is below the threshold are invisible to the observer. Then one can say that, in this region, the observer is invariant to the transform. Here we propose a methodology that allows us to assign specific invisibility thresholds per metric. This proposal uses transduction functions to a common internal representation based on a subjectively rated database and a psychophysically measured threshold in this representation.
Then, we can (1) assess if the thresholds for the metrics are comparable to those found for human observers, and (2) assess if the sensitivities of a metric for the different distortions follow the same order as the human sensitivities (for instance humans are more sensible to rotation changes than to illumination changes). This proposed evaluation of invariance to distortions is a necessary complement to the conventional evaluation of the visibility of distortions because, as we will see, none of the studied metrics presents human-like thresholds nor sensitivities for all the transformations considered.
The structure of the paper is as follows: Section II describes the proposed methodology to assess the human-like behavior of metric invariances. The proposed methodology consists of comparing the detection thresholds for humans and metrics. This proposal depends on several concepts (transduction functions, psychophysical thresholds for humans, theoretical thresholds for metrics… ) that will be detailed in this section too. Section III describes the experimental setting, and Section IV considers the results on the thresholds and the sensitivities. Finally, the discussion and conclusions are presented in Section V and Section VI, respectively.
II Proposed Methodology: Thresholds and Sensitivities
@@
While invariance thresholds in human observers, $\theta_{\tau}^{\textrm{H}}$, are easy to understand and measure [24, 25, 26, 27], they are not obvious to define in artificial models of perceptual distance. The reason is simple: for the usual image quality models any non-zero image distortion leads to non-zero variations in the distance. In this situation where artificial distances are real-valued functions one should define a value, the threshold distance, $\mathcal{D}_{\tau}$, below which the difference between the images could be disregarded (or taken as zero). Once this threshold distance is available (eventually in a common scale for all metrics) one can translate this threshold to the axis that measures the distortion intensity, $\theta$, and hence obtain the threshold of the metric, $\theta_{\tau}^{M}$, in the same units that has been measured for humans, $\theta_{\tau}^{H}$. See an illustration of this concept in Figure 1.
@@
Figure 1: Illustration of the methodology: comparison of metric vs human thresholds. For each distortion, parametrized with $\theta$, we evaluate three different metrics (continuous curves in blue, red, and green) in common units of subjective distance ($\mathcal{D}$, the normalized Mean Opinion Score). Given $N$ competing metrics, these curves are the transduction functions that map the physical description of the distortion intensity, $\theta$ into a common scale of subjective distance, $\mathcal{D}_{M}=g_{M}(\theta)$, with $M=1,\ldots,3$. The yellow lines show the human threshold in the internal distance representation, $\mathcal{D}_{\tau}$, experimentally measured in this work with a psychometric function. We assign the metric thresholds from the generic human threshold using the inverse of the transduction functions for the different metrics: $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. Given the fact that $\mathcal{D}_{\tau}$ has an experimental uncertainty, the inverse gives us an interval - dotted lines- on the x-axis where the human threshold should fall for each metric and this particular distortion. For any number $N$ of metrics we propose to compare the metric thresholds, $\theta_{\tau}^{M}$, for $M=1,\ldots,N$, with the actual human threshold in intensity units, $\theta_{\tau}^{H}$. In this case, $\theta_{\tau}^{H}$ (the black line), falls within the red interval. Therefore, the metric 2 is the one with the human-like behavior.
@@
The diagram in Figure 1 displays the relation between the physical description of the intensity of certain distortions in abscissas (the parameter $\theta$, e.g. the angle in a rotation transform), and the common internal description of the perceived distance in ordinates (the distance $\mathcal{D}$, e.g. the normalized Mean Opinion Score units explained below). This relation between the physical description of the intensity and common perceptual distance is what we call transduction function. The example displays three transduction functions, $\mathcal{D}_{M}=g_{M}(\theta)$, with $M=1,2,3$, in blue, red, and green, for the three corresponding metrics. Transduction functions are monotonic: the application of a progressively increasing distortion along the abscissas leads to monotonic increments in distances in ordinates. The threshold distance, $\mathcal{D}_{\tau}$, in the internal representation is plotted in orange in Figure 1. This threshold may be uncertain (as represented by the central value, solid line, and the quartile limits, in dashed lines), but the empirical transduction functions can be used to put this internal threshold back in the axis that describes the distortion intensity: $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. In this way, one can check if the actual invisibility threshold measured for humans (black line) is consistent with the threshold interval deduced for each metric. In our illustration, the metric 2 (red lines) is the only one compatible with human behavior.
@@
      A metric equalization function transforming the previous (non-scaled) distance values into the common scale of the internal distance representation (what we called normalized MOS units in the illustration). Here we propose to use auxiliary empirical data (e.g. certain subjectively rated databases) to scale the range of the different metrics: $\mathcal{D}_{M}=f_{M}(d_{M})$. This makes the different $\mathcal{D}_{M}$ comparable.
@@
II-A Transduction: response and equalization
@@
[…]
@@
where an exponential function with $a_{n}>0$ and $0<b_{n}\neq 1$ is chosen because of the nature of the DMOS, which changes rapidly for low distortion intensities, low values of $\theta$, and saturates for bigger distortions, big values of $\theta$. This is due to the maximum number of comparisons performed when building the database. An example of an equalization function is shown in Figure 2 as an example.
Figure 2: Example of equalization function fitted on TID13 data, for the metric PerceptNet, following Eq. 2.
@@
In our work, we compute these transduction functions for:
  • Six distance metrics (four state-of-the-art deep-learning metrics [3, 4, 32, 23] and two convenient references: the Euclidean metric, RMSE, and the classical SSIM [13]).
@@ II-B Human thresholds
B.1. Human thresholds in the common internal representation
To find out if the metrics behave similarly to humans we need to determine the human invisibility threshold, $\mathcal{D}_{\tau}$. This is the distance in normalized DMOS units from which humans can’t tell the difference between $i$ and $i^{\prime}$ for low distortion intensities $\theta$. For that, we need a database with ratings and opinions of observers, for which we can assign a threshold value of invisibility. In this case, we use the TID13 database. [31].
The value of $\mathcal{D}_{\tau}$ could be roughly estimated by visual inspection of the images presented in Appendix A: images with low values of normalized DMOS (below 0.3) cannot be discriminated from the original, while images with big normalized DMOS (above 0.6) are clearly distinct from the original. However, the more accurate estimation of such a threshold is obtained from the psychometric function in a constant stimulus experiment [30] applied to other fields within computer vision and image processing [36, 37, 38, 39, 40].
@@ B.1. Human thresholds in the common internal representation
In our experiment, the psychometric function has been evaluated in 20 different values of $\mathcal{D}$, repeated 15 times for 5 observers, resulting in $20\times 15\times 5=1500$ forced choices in total. The fitted psychometric function is shown in Figure 3 and shows that the value of the threshold for humans in the common internal distance representation is
@@
B.2. Human threshold in physical units
In principle, one could take the thresholds $\theta_{\tau}^{H}$ from the classical literature [24, 25, 26, 27] but, considering the critical effect of the stimuli used in the experiments and the fact that classical thresholds were measured with synthetic stimuli, we decided to measure these thresholds for natural images to ensure consistency with the kind of images used by the metrics. We measure the thresholds as in Section II-B.1., according to the constant stimuli setting to derive the classical psychometric function [30]. In particular, images from ImageNet have been used to find the invisibility thresholds in a more realistic context.
@@ B.2. Human threshold in physical units
Figure 4: Experimental (human) psychometric functions measured with natural images. According to them, the human thresholds for the geometric affine transformations are 0.23 $\pm$ 0.10 degrees for translation, 3.6 ±$\pm 1.5$ degrees for rotation, and 1.03 ±$\pm 0.02$ for scale.
On the other hand, now we address the results of the color thresholds which are two-dimensional in nature. As a convenient summary, here we describe the chromatic content of natural images using a single two-dimensional point in the so-called chromatic diagram [41]. In particular, we will use the chromatic coordinates of the spectrum of the illuminant assumed to be the illumination source of the scene. Moreover, we will describe the color-shifts introduced to check the invariance of metrics by the new chromatic coordinates of the new spectrum applied to the scene (as affine transform on the image vector as in Eq. 8 in Section III-A). Here we assume that the natural images in the databases are, in principle, illuminated by a generic white spectrum. Therefore, we describe the chromaticity of these images using the location of that generic white spectrum in the CIE xy diagram, i.e. in the 2d-point $(1/3,1/3)$ [41].
The two-dimensional nature of chromatic shifts implies that invisibility regions are not intervals $[0,\theta_{\tau}^{H}]$, but two-dimensional shapes (e.g. discrimination ellipses [27]). In order to determine the shape of such discrimination region around a certain central chromatic location, e.g. the aforementioned white point, one could measure one-dimensional intervals moving the images in different chromatic directions away from the selected center. For instance, Figure 5 represents the white point $(1/3,1/3)$, generic illuminant, and four linear paths indicating color shifts away from this center (radial gray lines). Measurements along radial lines can be done with the constant-stimulus + psychometric function paradigm mentioned above. In our case, given the ellipsoidal shape of the classical chromatic discrimination thresholds [27] (also shown in Figure 5), measuring only in four directions is enough for a trustable fit of an ellipse [42]. The fitted psychometric functions for these four chromatic directions are shown in Figure 6. The thresholds in departure in chromatic coordinates from these psychometric functions determine four two-dimensional experimental thresholds that allow to fit the ellipse in pink in Figure 5.
@@
Interestingly the ellipse we got by assessing the detection thresholds with color-shifted natural images is quite consistent with previous results that used synthetic stimuli and different experimental techniques [27, 43]. On the one hand, note the similarity in size and orientation between the pink and the purple ellipses. Note that the ellipse is wider in the yellow-blue (YB) direction than in the red-green (RG) direction. The authors in [43] reported that their threshold method with synthetic textured stimuli led to ellipses ×$\times 4.5$ than MacAdam ellipses [27] measured with the variance of color matching paradigms. As Fig 5 shows the MacAdam ellipses with a ×$\times 5$ visualization factor [41], our experimentally fitted pink ellipse obtained with thresholds from natural images is very consistent with [43].
@@
Table I summarizes the results with synthetic (obtained from classical literature) and real stimuli (experimentally measured in this work).
TABLE I: Human thresholds for the different related transformations. In the second column, we have the thresholds according to the classical literature. In the third column, we have the thresholds obtained in experiments with humans carried out in the laboratory.
@@
Translation 0.024 degrees [24] 0.23 $\pm$ 0.10 degrees
@@
Color discriminant MacAdam ellipse [27] Experimental ellipse (Figure 5)
From Table I we see that the rotation and scale thresholds are practically equivalent, but this is not the case for the translation threshold. This can be attributed to the fact their experiments with synthetic stimuli [24] can be considered as being close to a hyperacuity setting, (which returns very low thresholds) [44]. Our experiments employ natural images and do not enforce the hyperacuity setting. This results in a considerably higher detection threshold for this transformation. Recent work we could consider is the thresholds measured with gabor filters [45]. In this case, the human rotation threshold was considered to be 2.7 degrees of rotation and the human translation threshold the equivalent of 0.12 degrees of translation. These measurements are quite consistent with ours. Taking into account that the methodology of obtaining the threshold is the same as in our case, the difference in the thresholds is given exclusively by the type of stimulus used. For this reason, the detection threshold for translation from other literature will not be considered in the results because of differences in the procedure for obtaining it and its interpretation in the case of classical literature; and for the type of stimulus used in the recent literature.
II-C Metric thresholds in physical units
Following the methodology in Figure 1, we can obtain the metric thresholds in physical units, $\theta_{\tau}^{M}$. For example, the rotation threshold is expressed in degrees. Through the function $g_{M}(\theta)$, specific to each metric, a value in physical units is assigned to the threshold in the common internal representation, computed as $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. These values can be compared numerically with the thresholds obtained for humans, as they are expressed in the same units because they were obtained with the psychometric functions, Section 3. Therefore, a metric can be said to have human behavior if the invisibility threshold for a certain affine transformation coincides, $\theta_{\tau}^{H}=\theta_{\tau}^{M}$. In particular, since $\mathcal{D}_{\tau}$ has an associated uncertainty, $\theta_{\tau}^{M}$ also has a confidence interval. In this case, we will check whether the human threshold falls within the confidence interval of each metric.
This test is particularly demanding because it is strictly quantitative, i.e. numerical comparison. It is to be expected, that the metrics will not be able to reproduce these thresholds. This is why an alternative, less demanding, test is proposed, reflecting the qualitative behavior of the metrics in the face of these distortions.
II-D Sensitivities of metrics and humans for the different distortions

### v3

# Data Curation, Investigation, Software, Visualization

@@ Abstract
Subjective image quality metrics are usually evaluated according to the correlation with human opinion in databases with distortions that may appear in digital media. However, these oversee affine transformations which may represent better the changes in the images actually happening in natural conditions. Humans can be particularly invariant to these natural transformations, as opposed to the digital ones.
In this work, we propose a methodology to evaluate any image quality metric by assessing their invariance to affine transformations, specifically: rotation, translation, scaling, and changes in spectral illumination. Here, invariance refers to the fact that certain distances should be neglected if their values are below a threshold. This is what we call invisibility threshold of a metric. Our methodology consists of two elements: (1) the determination of a visibility threshold in a subjective representation common to every metric, and (2) a transduction from the distance values of the metric and this common representation. This common representation is based on subjective ratings of readily available image quality databases. We determine the threshold in such common representation (the first element) using accurate psychophysics. Then, the transduction (the second element) can be trivially fitted for any metric: with the provided threshold extension of the method to any metric is straightforward. We test our methodology with some well-established metrics and find that none of them show human-like invisibility thresholds.
[…]
keywords:
[…]
[1]https://github.com/Rietta5/InvarianceTestIQA.
{graphicalabstract}[Uncaptioned image]
{highlights}
We propose a methodology to assign invisibility thresholds for any perceptual metric.
We measure the human invisibility threshold to image distortions in the Difference Mean Opinion Score representation by fitting a psychometric function.
We expand classical literature on human thresholds to affine transformations in physical units using natural stimuli as opposed to synthetic stimuli.
The proposed methodology takes the above-mentioned thresholds and straightforward fitting of transduction functions to derive thresholds for any metric in physical units.
Alignment between metrics and humans is assessed in two ways: (1) a quantitative comparison of thresholds, and (2) a qualitative comparison of sensitivities.
1 Introduction
Artificial models used to predict subjective image quality, usually referred to as perceptual metrics, are often used as measures to optimize other image processing tasks [8, 9, 5, 42] so, assessing their performance is critical. Usually, these models are evaluated (or even tuned) according to their ability to correlate with human opinion in databases including a wide range of generic distortions [13, 25, 48]. In fundamental terms, this means trying to predict the visibility of generic distortions. However, focus on generic distortions may be a problem to take into account other relevant phenomena [30], and human vision may also be described in terms of perceptual constancies or invariances [10, 12, 17, 28, 32, 46].
@@ 1 Introduction
In fact, the spirit of the influential SSIM was focused on measuring changes of structure [43] to achieve invariances to irrelevant transformations [44]. Moreover, Wang and Simoncelli [45] decomposed generic distortions into structural and non-structural components so that the part not affecting the structure (e.g. affine transformations) could be processed, and weighted, differently. On the other hand, metrics with bio-inspired, explainable architecture [13, 20, 21, 22, 27, 29, 30, 31] work in multi-scale / multi-orientation representations where invariances can be introduced by means of appropriate poolings in the representations [3]. This sort of poolings are thought to happen in the visual brain leading to invariances and texture metamers [11].
@@
In this work, we propose a methodology to evaluate the metrics from the point of view of human detection thresholds: by assessing the invariance of the metrics to image transformations that are irrelevant (or invisible) to human observers. In particular, we compare the ability of metrics to be invariant to affine transformations in the same way as humans are. For example, the classical literature on visual thresholds determines the intensity of certain affine transformations which is invisible for humans [24, 26, 36, 40]. The sizes of invisibility thresholds are related to the more general concept of invariance to transformations. By definition, transformations whose intensity is below the threshold are invisible to the observer. Then one can say that, in this region, the observer is invariant to the transform. Here, we propose a methodology that allows us to measure the invisibility thresholds for any given metric. The proposed methodology has two elements: (1) the determination of a visibility threshold in a subjective representation common to every metric, and (2) a transduction function from the raw distance values of the metric and this common representation. This common representation is based on subjective ratings of readily available image quality databases. Here we present accurate psychophysical measurements of the threshold in such subjective representation (the first element), and then, the transduction function (second element) can be trivially fitted for any metric using available image quality databases. In this way the proposed methodology is general and applicable to any metric any metric.
Then, we can (1) assess if the thresholds for the metrics are comparable to those found for human observers, and (2) assess if the sensitivities of a metric for different distortions follow the same order as the human sensitivities (for instance, humans are more sensible to rotation changes than to illumination changes). This proposed evaluation of invariance to distortions is a necessary complement to the conventional evaluation of the visibility of distortions because, as we will see, none of the studied metrics presents human-like thresholds nor sensitivities for all the transformations considered.
The structure of the paper is as follows: Section 2 describes the proposed methodology to assess the human-like behavior of metric invariances, consisting of comparing the detection thresholds for humans and metrics. This proposal depends on several concepts (transduction functions, psychophysical thresholds for humans, theoretical thresholds for metrics… ) that will be detailed in this section too. Section 3 describes the experimental setting, and Section 4 considers the results on the thresholds and the sensitivities. The discussion and conclusions are presented in Section 5 and Section 6, respectively.
2 Proposed Methodology: Thresholds and Sensitivities
@@
While invariance thresholds in human observers, $\theta_{\tau}^{\textrm{H}}$, are easy to understand and measure [24, 26, 36, 40], they are not obvious to define in artificial models of perceptual distance. The reason is simple: for the usual image quality models, any non-zero image distortion leads to non-zero variations in the distance. In this situation, where artificial distances are real-valued functions, one should define a value, the threshold distance, $\mathcal{D}_{\tau}$, below which the difference between the images could be disregarded (or taken as zero). Once this threshold distance is available (eventually in a common scale for all metrics) one can translate this threshold to the axis that measures the distortion intensity, $\theta$, and hence obtain the threshold of the metric, $\theta_{\tau}^{M}$, in the same units that has been measured for humans, $\theta_{\tau}^{H}$. See an illustration of this concept in Figure 1.
@@
Figure 1: Illustration of the methodology: comparison of metric vs human thresholds. For each distortion, parametrized with $\theta$, we evaluate three different metrics (continuous curves in blue, red, and green) in common units of subjective distance ($\mathcal{D}$, the normalized Differential Mean Opinion Score). Given NN competing metrics, these curves are the transduction functions that map the physical description of the distortion intensity, $\theta$ into a common scale of subjective distance, $\mathcal{D}_{M}=g_{M}(\theta)$, here with $M=1,\ldots,3$. The yellow lines show the human threshold in the internal distance representation, $\mathcal{D}_{\tau}$, experimentally measured in this work with a psychometric function. We assign the metric thresholds from the generic human threshold using the inverse of the transduction functions for the different metrics: $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. Given the fact that $\mathcal{D}_{\tau}$ has an experimental uncertainty, the inverse gives us an interval - dotted lines- on the x-axis where the human threshold should fall for each metric and this particular distortion. For any number $N$ of metrics we propose to compare the metric thresholds, $\theta_{\tau}^{M}$, for $M=1,\ldots,N$, with the actual human threshold in intensity units, $\theta_{\tau}^{H}$. In this case, $\theta_{\tau}^{H}$ (the black line), falls within the red interval. Therefore, the metric 2 is the one with the human-like behavior.
@@
The diagram in Figure 1 displays the relation between the physical description of the intensity of certain distortions in abscissas (the parameter $\theta$, e.g. the angle in a rotation transform), and the common internal description of the perceived distance in ordinates (the distance $\mathcal{D}$, e.g. the normalized Differential Mean Opinion Score units explained below). This relation between the physical description of the intensity and common perceptual distance is what we call transduction function. The example displays three transduction functions, $\mathcal{D}_{M}=g_{M}(\theta)$, with $M=1,2,3$, in blue, red, and green, for the three corresponding metrics. Transduction functions are monotonic: the application of a progressively increasing distortion along the abscissas leads to monotonic increments in distances in ordinates. The threshold distance, $\mathcal{D}_{\tau}$, in the internal representation is plotted in orange in Figure 1. This threshold may be uncertain (as represented by the central value, solid line, and the quartile limits, in dashed lines), but the empirical transduction functions can be used to put this internal threshold back in the axis that describes the distortion intensity: $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. In this way, one can check if the actual invisibility threshold measured for humans (black line) is consistent with the threshold interval deduced for each metric. In our illustration, the metric 2 (red lines) is the only one compatible with human behavior.
@@
      A metric equalization function transforming the previous (non-scaled) distance values into the common scale of the internal distance representation (what we called normalized DMOS units in the illustration). Here we propose to use auxiliary empirical data (e.g. certain subjectively rated databases) to scale the range of the different metrics: $\mathcal{D}_{M}=f_{M}(d_{M})$. This makes the different $\mathcal{D}_{M}$ comparable.
@@
2.1 Transduction: response and equalization
@@
In this work, we propose to use auxiliary empirical data to determine a common distance scale, $\mathcal{D}$, for any metric. In particular, subjectively rated image quality databases (e.g. TID [35]) consist of pairs of original and distorted images, $\{(i^{p},{i^{p}}^{\prime})\}_{p=1}^{P}$, with associated subjective scores, the so-called Differential Mean Opinion Scores, $\{\textrm{DMOS}^{p}\}_{p=1}^{P}$. The databases contain a wide set of generic distortions of different intensities thus ranging from invisible distortions to highly noticeable distortions. In this setting, DMOS values in the database can be normalized to be in the [0,1] range. In this way, the lower and higher values of the normalized DMOS represent an invisible distortion and the biggest subjective distortion in the database respectively. Therefore, if the range of distortions in the database is wide, the variations induced by $T_{\theta}$ will be within the limits of the normalized DMOS, and hence, it can be used to set the common distance scale $\mathcal{D}=\textrm{norm DMOS}\in[0,1]$.
@@
where an exponential function with $a_{n}>0$ and $0<b_{n}\neq 1$ is chosen because of the nature of the DMOS, which changes rapidly for low distortion intensities, low values of $\theta$, and saturates for bigger distortions, big values of $\theta$. An example of an equalization function is shown in Figure 2 as an example.
Figure 2: Example of equalization function fitted on TID13 data, for the metric LPIPS, following Eq. 2.
@@
In our work, as means of showcasing our proposed methodology, we compute these transduction functions for:
  • Six distance metrics: four state-of-the-art deep-learning metrics [2, 7, 13, 48], and two convenient references (the Euclidean metric, RMSE, and the classical SSIM [43]).
@@ 2.2 Human thresholds
2.2.1 Human thresholds in the common internal representation
$\mathcal{D}_{\tau}$ is the distance in normalized DMOS units from which humans can’t tell the difference between $i$ and $i^{\prime}$ for low distortion intensities $\theta$. For that, we need a database with ratings and opinions of observers, for which we can assign a threshold value of invisibility. In this case, we use the TID13 database [35].
The value of $\mathcal{D}_{\tau}$ could be roughly estimated by visual inspection of the images presented in Appendix A (example in Figure 3): images with low values of normalized DMOS (below 0.3) cannot be discriminated from the original, while images with big normalized DMOS (above 0.6) are clearly distinct from the original. However, the more accurate estimation of such a threshold is obtained from the psychometric function in a constant stimulus experiment [18] applied to other fields within computer vision and image processing [38, 37, 16, 41, 33].
Figure 3: Example image from Appendix A, showing the relationship between the DMOS of the images and their visibility. An image with a low DMOS (around 0.3) is indistinguishable from the original, while a high DMOS (around 0.6) is clearly different.
@@ 2.2.1 Human thresholds in the common internal representation
In our experiment, the psychometric function has been evaluated in 20 different values of $\mathcal{D}$, repeated 15 times for 11 observers, resulting in $20\times 15\times 11=3300$ forced choices in total. The fitted psychometric function is shown in Figure 4 and shows that the value of the threshold for humans in the common internal distance representation is:
@@
2.2.2 Human threshold in physical units
In principle, one could take the thresholds $\theta_{\tau}^{H}$ from the classical literature [24, 26, 36, 40] but, considering the critical effect of the stimuli used in the experiments and the fact that classical thresholds were measured with synthetic stimuli, we decided to measure these thresholds for natural images to ensure consistency with the kind of images used by the metrics. We measure the thresholds according to the constant stimuli setting, as in Section 2.2.1, to derive the classical psychometric function [18]. In particular, images from ImageNet have been used to find the invisibility thresholds in a more realistic context.
@@ 2.2.2 Human threshold in physical units
Figure 5: Experimental (human) psychometric functions measured with natural images. According to them, the human thresholds for the geometric affine transformations are 0.23 $\pm$ 0.10 degrees for translation, 3.6 $\pm 1.5$ degrees for rotation, and 1.03 $\pm 0.02$ for scale.
On the other hand, now we address the results of the color thresholds which are two-dimensional in nature. As a convenient summary, here we describe the chromatic content of natural images using a single two-dimensional point in the so-called chromatic diagram [47]. In particular, we will use the chromatic coordinates of the spectrum of the illuminant assumed to be the illumination source of the scene. Moreover, we will describe the color-shifts introduced to check the invariance of metrics by the new chromatic coordinates of the new spectrum applied to the scene (as affine transform on the image vector as in Eq. 8 in Section 3.1). Here we assume that the natural images in the databases are, in principle, illuminated by a generic white spectrum. Therefore, we describe the chromaticity of these images using the location of that generic white spectrum in the CIE xy diagram, i.e. in the 2d-point $(1/3,1/3)$ [47].
The two-dimensional nature of chromatic shifts implies that invisibility regions are not intervals, but two-dimensional shapes (e.g. discrimination ellipses [26]). In order to determine the shape of such discrimination region around a certain central chromatic location, e.g. the aforementioned white point, one could measure one-dimensional intervals moving the images in different chromatic directions away from the selected center. For instance, Figure 6 represents the white point $(1/3,1/3)$, generic illuminant, and four linear paths indicating color shifts away from this center (radial gray lines). Measurements along radial lines can be done with the constant-stimulus paradigm mentioned above. In our case, given the ellipsoidal shape of the classical chromatic discrimination thresholds [26] (also shown in Figure 6), measuring only in four directions is enough for a trustable fit of an ellipse [34]. The fitted psychometric functions for these four chromatic directions are shown in Figure 7. The thresholds in departure in chromatic coordinates from these psychometric functions determine four two-dimensional experimental thresholds that allow to fit the ellipse in pink in Figure 6.
@@
Interestingly the ellipse we got by assessing the detection thresholds with color-shifted natural images is quite consistent with previous results that used synthetic stimuli and different experimental techniques [26, 15]. On the one hand, note the similarity in size and orientation between the pink and the purple ellipses. Note that the ellipse is wider in the yellow-blue (YB) direction than in the red-green (RG) direction. The authors in [15] reported that their threshold method with synthetic textured stimuli led to ellipses $\times 4.5$ than MacAdam ellipses [26] measured with the variance of color matching paradigms. As Fig 6 shows the MacAdam ellipses with a $\times 5$ visualization factor [47], our experimentally fitted pink ellipse obtained with thresholds from natural images is very consistent with [15].
@@
Table 1 summarizes the results with synthetic (obtained from classical literature) and real stimuli (experimentally measured in this work).
Table 1: Human thresholds for the different related transformations. In the second column, we have the thresholds according to the classical literature. In the third column, we have the thresholds obtained in experiments with humans carried out in the laboratory.
@@
Translation
0.024 degrees [24]
0.12 degrees [1]
0.23 $\pm$ 0.10 degrees
@@
Color discrimination MacAdam ellipse [26] Experimental ellipse (Fig. 6)
From Table 1 we see that the rotation and scale thresholds are practically equivalent, but this is not the case for the translation threshold. This can be attributed to the fact their experiments with synthetic stimuli [24] can be considered as being close to a hyperacuity setting, (which returns very low thresholds) [39]. Our experiments employ natural images and do not enforce the hyperacuity setting. This results in a considerably higher detection threshold for this transformation. Recent work we could consider is the thresholds measured with gabor filters [1]. In this case, the human rotation threshold was considered to be 2.7 degrees of rotation and the human translation threshold the equivalent of 0.12 degrees of translation. These measurements are quite consistent with ours. Taking into account that the methodology of obtaining the threshold is the same as in our case, the difference in the thresholds is given exclusively by the type of stimulus used. For this reason, the detection threshold for translation from other literature will not be considered in the results because of differences in the procedure for obtaining it and its interpretation in the case of classical literature; and for the type of stimulus used in the recent literature.
2.3 Metric thresholds in physical units
Following the methodology in Figure 1, we can obtain the metric thresholds in physical units, $\theta_{\tau}^{M}$. For example, the rotation threshold is expressed in degrees. Through the function $g_{M}(\theta)$, specific to each metric, a value in physical units is assigned to the threshold in the common internal representation, computed as $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. These values can be compared numerically with the thresholds obtained for humans, as they are expressed in the same units because they were obtained with the psychometric functions, Section 2.2.2. Therefore, a metric can be said to have human behavior if the invisibility threshold for a certain affine transformation coincides, $\theta_{\tau}^{M}=\theta_{\tau}^{H}$. In particular, since $\mathcal{D}_{\tau}$ has an associated uncertainty, $\theta_{\tau}^{M}$ also has a confidence interval. In this case, we will check whether the human threshold falls within the confidence interval of each metric.
This test is particularly demanding because it is strictly quantitative. It is to be expected, that the metrics will not be able to reproduce these thresholds. This is why an alternative, less demanding, test is proposed, reflecting the qualitative behavior of the metrics in the face of these distortions.
2.4 Sensitivities of metrics and humans for the different distortions